OpenStudy (anonymous):

Find the value of cos^2 15 exactly

OpenStudy (luigi0210):

\[\cos^2 15\] That?

OpenStudy (anonymous):

yea

OpenStudy (callisto):

Let A = 15 cos A = cos 15 cos (2A) = cos (2x15) = cos 30. Do you know what cos 30 is?

OpenStudy (anonymous):

sqrt(3)/2

OpenStudy (anonymous):

then what??...

OpenStudy (callisto):

Use the identity: \[cos2A = 2cos^2A -1\]to get \(cos^2A\)

OpenStudy (anonymous):

so is the answer sqrt(3) + 2?

OpenStudy (callisto):

Maybe you'd like to check your calculations again.

OpenStudy (anonymous):

ok can you do it for me?

OpenStudy (callisto):

No, but I don't mind checking it for you.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so you go:\[\frac{ \sqrt{3} }{ 2 } + 1 = 2 \cos^2 \]

OpenStudy (anonymous):

then\[\frac{ (\frac{ \sqrt{3} }{ 2 } + 1) }{ 2 } = \cos^2 A\]

OpenStudy (anonymous):

then what after that?

OpenStudy (callisto):

Almost there. Simplify \(\frac{ (\frac{ \sqrt{3} }{ 2 } + 1) }{ 2 }\)

OpenStudy (anonymous):

so sqrt(3) + 1?

OpenStudy (callisto):

No.

OpenStudy (anonymous):

sqrt(3) + 2?

OpenStudy (callisto):

No...

OpenStudy (anonymous):

then how

OpenStudy (callisto):

\[\frac{ (\frac{ \sqrt{3} }{ 2 } + 1) }{ 2 } = (\frac{ \sqrt{3} }{ 2\times2 } + \frac{1}{2}) \]You may want to further simplify it.